A women is 7 times as old as her daughter. If in five years time she is four times as old as her daughter, what are their present ages?
Answer by mananth(16946)

daughter =x
woman = 7x
5 years later
daughter = x+5
woman = 7x+5
...
7x+5 = 4(x+5)
7x+5=4x+20
7x-4x=20-5
3x=15
x= 5 years daughter's age now
Woman's age = 7*5 = 35 years now